the width of a rectangle is 4 meters less than twice its length. If its area is 70 square meters, find the dimensions of the rectangle. Please show the formula/equation.
Answer by checkley77(12844)

W=2L-4
WL=70
(2L-4)L=70
2L^2-4L-70=0
2(L^2-2L-35)=0
2(L-7)(L+5)=0
L-7=0
L=7 IS THE LENGTH OF THE RECTANGLE.
W=2*7-4
W=14-4
W=10 IS THE WIDTH OF THE RECTANGLE.
PROOF:
7*10=70
70=70